# On anomalous dissipation induced by transport noise

### v2

@@ 4. Anomalous dissipation of enstrophy – 2D NSEs
This section is devoted to the proof of Theorem 1.1. Its proof follows the one of Theorem 1.2 given in Section 3 as the convective nonlinearity creates only minor difficulties. This section is organized as follows. In Subsection 4.1 we first show the global well-posedness of the 2D NSEs with cut-off, where the cut-off is used to tame the nonlinearity. Secondly, in Subsection 4.2, we prove Theorem 1.1 by removing the cut-off and arguing as in the case of passive scalars.
@@ 4.1. Scaling limit with cut-off at a fixed Reynolds number
(4.1) $\partial_{t}\zeta_{{\rm cut}}^{\nu}&+\phi_{R,r}(\zeta_{{\rm cut}}^{\nu})\,[({\mathcal{K}}\,\zeta_{{\rm cut}}^{\nu}\cdot\nabla)\zeta_{{\rm cut}}^{\nu}]=\nu\Delta\zeta_{{\rm cut}}^{\nu}\\ &\qquad\qquad\qquad+\sqrt{2\mu}\sum_{k\in{\mathbb{Z}}^{2}_{0}}\theta_{k}\,(\sigma_{k}\cdot\nabla)\zeta_{{\rm cut}}^{\nu}\circ\dot{W}_{t}^{k}&\text{ on }&\mathbb{T}^{2},\\ \zeta_{{\rm cut}}^{\nu}(0,\cdot)&=\zeta_{0}&\text{ on }&\mathbb{T}^{2},\end{aligned}\right.$
where
@@
Other possible choices of the norm in the cut-off function are possible for the 2D NSEs in vorticity formulation (1.1). However, the above cut-off has the highest smoothness that is compatible with Meyers’ estimate in Theorem A.1. Moreover, the choice of $H^{r}$ with $r>0$ will be crucial in Section 5 below, where we deal with 2D NSEs in velocity formulation which are critical in $L^{2}$.
@@
Let $p\in[2,\infty)$. Fix $\nu>0$, $\theta^{\nu}\in\ell^{2}$ and $\zeta_{0}\in B^{1-2/p}_{2,p}(\mathbb{T}^{2})$. Let $\tau^{\nu}:\Omega\to[0,\infty]$ and $\zeta_{{\rm cut}}^{\nu}:[0,\infty)\times\Omega\to H^{1}(\mathbb{T}^{2})$ be a stopping time and a progressive measurable process, respectively.
  • We say that $(\zeta_{{\rm cut}}^{\nu},\tau^{\nu})$ is a local $p$-solution to (4.1) if the following hold:
    • $\zeta_{{\rm cut}}^{\nu}\in L^{p}_{{\rm loc}}([0,\infty);H^{1}(\mathbb{T}^{2}))\cap C([0,\infty);B^{1-2/p}_{2,p}(\mathbb{T}^{2}))$ a.s.;
    • ${\mathcal{K}}\,\zeta_{{\rm cut}}^{\nu}\zeta_{{\rm cut}}^{\nu}\in L^{p}_{{\rm loc}}([0,\infty);L^{2}(\mathbb{T}^{2};{\mathbb{R}}^{2}))$ a.s.;
    • a.s. for all $t\in{\mathbb{R}}_{+}$ it holds that
@@
      $+\sqrt{2\mu}\int_{0}^{t}\mathbf{1}_{[0,\tau^{\nu})}\Big{(}(\theta^{\nu}_{k}\,\sigma_{k}\cdot\nabla)\zeta_{{\rm cut}}^{\nu}(s)\Big{)}_{k}\,\mathrm{d}\mathcal{W}_{\ell^{2}}.$
  • A local $p$-solution $(\zeta_{{\rm cut}}^{\nu},\tau^{\nu})$ to (4.1) is said to be a unique local $p$-solution to (4.1) if for any local solution $(\xi_{{\rm cut}}^{\nu},\lambda^{\nu})$ we have $\lambda^{\nu}\leq\tau^{\nu}$ a.s. and $\xi_{{\rm cut}}^{\nu}=\zeta_{{\rm cut}}^{\nu}$ a.e. on $[0,\lambda^{\nu})\times\Omega$.
  • A unique local $p$-solution $(\zeta_{{\rm cut}}^{\nu},\tau^{\nu})$ to (4.1) is said to be a unique global $p$-solution to (4.1) if $\tau^{\nu}=\infty$ a.s.
For the optimality of the regularity of the initial data, the reader is referred to the comments below Theorem A.1. Similar to Definition 2.2, if $p=2$, then we simply write ‘solution’ instead of ‘$p$-solution’.
@@
Fix $\nu\in(0,1)$ and $\mu>0$. Then there exists $p_{0}(\nu,\mu)>2$ for which the following assertion holds. If for $p\in(2,p_{0}]$, $r\in(0,1-2/p)$ and $R>0$ we have
@@
    there exists a unique solution $\zeta_{{\rm det}}^{\nu}\in C([0,1];H^{r})\cap L^{2}(0,1;H^{1})$ on $[0,1]$ to
    (4.2) $\partial_{t}\zeta_{{\rm det}}^{\nu}+\phi_{R,r}(\zeta_{{\rm det}})\,[({\mathcal{K}}\,\zeta_{{\rm det}}^{\nu}\cdot\nabla)\zeta_{{\rm det}}^{\nu}]&=(\nu+\mu)\Delta\zeta_{{\rm det}}^{\nu}&\text{ on }&\mathbb{T}^{2},\\ \zeta_{{\rm det}}^{\nu}(0,\cdot)&=u_{0}&\text{ on }&\mathbb{T}^{2};\end{aligned}\right.$
@@
(4.3) $\lim_{n\to\infty}\mathbf{P}\Big{(}\sup_{t\in[0,1]}\|\zeta_{{\rm cuts},n}^{\nu}(t)-\zeta_{{\rm det}}^{\nu}(t)\|_{H^{r_{0}}}\geq\varepsilon\Big{)}=0\ \text{ for all }\varepsilon>0,$
where $\zeta_{{\rm cuts},n}^{\nu}$ is the unique global $p$-solution to (4.1) with $\theta=\theta^{n}$ and initial data $\zeta_{0,n}$.
As for Proposition 3.1, the key point is that the Sobolev space in (4.3) has positive smoothness. Moreover, repeating the argument in Remark 3.4, there is no convergence of gradients of $\zeta_{{\rm cuts},n}^{\nu}$ to $\zeta_{{\rm det}}^{\nu}$ in $L^{2}(0,1;L^{2})$. Moreover, in contrast to the linear case, the uniqueness of (weak) solutions to (4.2) is not immediate and therefore we include it as an assumption.
The existence of a unique global $p$-solution to (4.1) is proven in Lemma 4.3 below. Unique local solutions to (4.2) on $[0,1]$ are as in Definition 4.1 with trivial noise, $p=2$ and life-time $\geq 1$. Note that the condition $\zeta_{{\rm det}}^{\nu}\in C([0,1];H^{r})$ is not a-priori given in the requirements for being local solutions, this is why it has been displayed in (2). Similarly, $\zeta_{{\rm det}}^{\nu}\in L^{2}(0,1;H^{1})$ is needed as life-time $\geq 1$ does not imply that $t\mapsto\|\zeta_{{\rm det}}^{\nu}(t)\|_{L^{2}}^{2}$ is integrable near $t=1$.
The proof of Proposition 4.2 is given at the end of this subsection. Following the argument of Subsection 3.1, the key ingredients are the following estimates which are uniform in the class of normalized radially symmetric $\theta^{\nu}$.
@@
As in the proof of Lemma 3.2, the key point is the independence of $(p_{0},C_{0})$ on $\theta$. As in the latter result, the key tools are the stochastic Meyers’ estimates of Theorem A.1. To handle the nonlinearity, we use the sub-criticality of $H^{r}$ for all $r>0$ in the case of 2D NSEs both in vorticity and velocity form (see Section 5 below for the latter). The sub-criticality of $H^{r}$ is exploited via the following inequality (cf., [1, Lemmas 4.3 and 4.5]): For all $r<\frac{1}{2}$,
(4.5) $\big{\|}\mathbb{P}[\nabla\cdot([{\mathcal{K}}\,\zeta]\,\zeta)]\big{\|}_{H^{-1}}^{2}\lesssim\|\zeta\|_{L^{4}}^{2}\lesssim\|\zeta\|_{H^{1/2}}^{2}\lesssim\|\zeta\|_{H^{r}}^{1+\kappa_{r}}\|\zeta\|_{H^{1}}^{1-\kappa_{r}}$
where $\kappa_{r}=\frac{r}{1-r}>0$. For later use (see Section 5 below), let us note that the (4.5) also holds if $[{\mathcal{K}}\,\zeta]\zeta$ is replaced by $v\otimes v$ with $v\in H^{1}$.
The above follows from the Sobolev embedding $H^{1/2}(\mathbb{T}^{2})\hookrightarrow L^{4}(\mathbb{T}^{2})$ as well as standard interpolation inequality. The sub-criticality of $H^{r}$ is encoded in the power $11-\kappa_{r}<$ 1 for the $H^{1}$-norm on the RHS(4.5). This fact will allow us to absorb the corresponding term via Young’s inequality.
Proof of Proposition 4.3.
@@
The proof is now split into three steps. We first prove the existence of a unique local $p$-solution $(\zeta_{{\rm cut}},\tau)$ with a corresponding blow-up criterion. Secondly, we prove that such a localized version of (4.4) for such a local solution. In the third step, we prove that $\tau=\infty$ a.s. (i.e., $(\zeta_{{\rm cut}},\tau)$ is global) and that (4.4) holds.
Step 1: There exists a unique local $p$-solution $(\zeta_{{\rm cut}},\tau)$ to (4.1) and for all $T\in(0,\infty)$ following blow-up criterion hold
@@
To prove the existence of a unique local $p$-solution to (4.1), we employ [4, Theorem 4.8]. To this end, we rewrite (4.1) as a stochastic evolution equation on $X_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}H^{-1}$:
@@
where, for $\zeta_{{\rm cut}}\in X_{1}\stackrel{{\scriptstyle{\rm def}}}{{=}}H^{1}$,
$A\zeta_{{\rm cut}}=-(\nu+\mu)\Delta\zeta_{{\rm cut}}$ $,\qquad F(\zeta_{{\rm cut}})=\phi_{R,r}(\zeta_{{\rm cut}})\mathbb{P}[\nabla([{\mathcal{K}}\,\zeta_{{\rm cut}}]\,\zeta_{{\rm cut}})],$
@@
Comparing Definition 4.1 with [4, Definition 4.4], one readily see that unique $p$-solution to (4.1) are equivalent to maximal $L^{p}_{0}$-solution to (4.7) (see also [5, Remark 5.6]). Set $X_{\beta}\stackrel{{\scriptstyle{\rm def}}}{{=}}H^{-1+2\beta}$ for $\beta\in(0,1)$. By [4, Theorem 4.8], it remains to show the existence of $1/2<\beta\leq\varphi\leq 1$ such that $\varphi+\beta<2-1/p$ such that, for all $n\geq 1$ and for all $\zeta,\zeta^{\prime}\in X_{1}$ such that $\|\zeta\|_{B^{1-2/p}_{2,p}},\|\zeta^{\prime}\|_{B^{1-2/p}_{2,p}}\leq n$,
(4.8) $\|F(\zeta)-F(\zeta^{\prime})\|_{X_{0}}$ $\lesssim_{n}(\|\zeta\|_{X_{3/4}}+\|\zeta^{\prime}\|_{X_{3/4}})\|\zeta-\zeta^{\prime}\|_{X_{3/4}}$
@@
To prove (4.8), let us write $F(\zeta)-F(\zeta^{\prime})=I_{\zeta,\zeta^{\prime}}+J_{\zeta,\zeta^{\prime}}$ where
$I_{\zeta,\zeta^{\prime}}$ $=\phi_{R,r}(\zeta)\,[\nabla\cdot([{\mathcal{K}}\,\zeta]\,\zeta)-\nabla\cdot([{\mathcal{K}}\,\zeta^{\prime}]\zeta^{\prime})],$
$J_{\zeta,\zeta^{\prime}}$ $=[\nabla\cdot([{\mathcal{K}}\,\zeta^{\prime}]\,\zeta^{\prime})]\big{(}\phi_{R,r}(\zeta)-\phi_{R,r}(\zeta^{\prime})\big{)}.$
The first term on the RHS(4.8) clearly follows by estimating $I_{\zeta,\zeta^{\prime}}$ and using $X_{3/4}=H^{1/2}(\mathbb{T}^{2})\hookrightarrow L^{4}(\mathbb{T}^{2})$. By arguing as in (4.5), one can check that, for all $r<\frac{1}{2}$,
$\big{\|}\mathbb{P}[\nabla\cdot([{\mathcal{K}}\,\zeta^{\prime}]\,\zeta^{\prime})]\big{\|}_{H^{-1}}^{2}\lesssim\|\zeta^{\prime}\|_{H^{r}}^{1+\widetilde{\kappa}_{r}}\|\zeta^{\prime}\|_{H^{1-r/2}}^{1+\widetilde{\kappa}_{r}}$
where $\widetilde{\kappa}_{r}\stackrel{{\scriptstyle{\rm def}}}{{=}}r/(2-3r)$. Since $B^{1-2/p}_{2,p}\hookrightarrow H^{r}$ as $1-2/p>r$ by assumption, the above inequality shows that the second term on the RHS(4.8) readily follows from estimating $J_{\zeta,\zeta^{\prime}}$ with $\varphi=1-\frac{r}{4}$ and $\beta=1-\frac{2}{p}+\varepsilon$ where $\varepsilon>0$ is arbitrary.
Finally, (4.6) follows from [5, Theorem 4.10(3)] applied to (4.7).
@@
where $\inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}\tau\wedge T$. The Meyers’ estimates of Theorem A.1 and a standard localization argument (see e.g., [4, Proposition 3.12(b)]) ensure that existence of $C_{1}(\nu,\mu)>0$ such that, for all $m\geq 1$, $p\in[2,p_{0}]$ and $u_{0}\in B^{1-2/p}_{2,p}$,
@@
Hence, $\|\zeta_{{\rm cut}}(t)\|_{B^{1-2/p}_{2,p}}^{p}+\mathbf{E}\int_{0}^{\tau\wedge T}\|\zeta_{{\rm cut}}(s)\|^{p}_{H^{1}}\,\mathrm{d}s<\infty$ a.s. on $\{\tau<T\}$, and therefore $\tau\geq T$ a.s. by (4.6). The arbitrariness of $T$ implies $\tau=\infty$ a.s., as desired.
Step 3: Proof of (4.4). The estimate (4.4) follows from (4.9) and the fact that $\tau=\infty$ a.s. by Step 1. ∎
@@
and that for all $\|u_{0}\|_{\mathbb{H}^{\delta}}\leq N$ the unique global solution $\zeta^{\nu}$ to (1.1) satisfies
@@
(4.10) $\partial_{t}\zeta_{{\rm det}}^{\nu}+({\mathcal{K}}\,\zeta_{{\rm det}}^{\nu}\cdot\nabla)\zeta_{{\rm det}}^{\nu}&=(\nu+\mu)\Delta\zeta_{{\rm det}}^{\nu}&\text{ on }&\mathbb{T}^{2},\\ \zeta_{{\rm det}}^{\nu}(0,\cdot)&=\zeta_{0}&\text{ on }&\mathbb{T}^{2}.\end{aligned}\right.$

### v3

@@ 4. Anomalous dissipation of enstrophy – 2D NSEs
This section is devoted to the proof of Theorem 1.1. Its proof follows the one of Theorem 1.2 given in Section 3 as the convective nonlinearity creates only minor difficulties. This section is organized as follows. In Subsection 4.1, we first show the global well-posedness of the 2D NSEs with cut-off, where the cut-off is used to tame the nonlinearity. Secondly, in Subsection 4.2, we prove Theorem 1.1 by removing the cut-off and arguing as in the case of passive scalars.
@@ 4.1. Scaling limit with cut-off at a fixed Reynolds number
(4.1) $\left\{\begin{aligned} \partial_{t}\zeta_{{\rm cut}}^{\nu}&+\phi_{R,r}(\zeta_{{\rm cut}}^{\nu})\,\big[({\mathcal{K}}\,\zeta_{{\rm cut}}^{\nu}\cdot\nabla)\zeta_{{\rm cut}}^{\nu}\big]=\nu\Delta\zeta_{{\rm cut}}^{\nu}\\ &\qquad\qquad\qquad+\sqrt{2\mu}\sum_{k\in{\mathbb{Z}}^{2}_{0}}\theta_{k}\,(\sigma_{k}\cdot\nabla)\zeta_{{\rm cut}}^{\nu}\circ\dot{W}_{t}^{k}&\text{ on }&\mathbb{T}^{2},\\ \zeta_{{\rm cut}}^{\nu}(0,\cdot)&=\zeta_{0}&\text{ on }&\mathbb{T}^{2},\end{aligned}\right$.
where $\theta=(\theta_{k})_{k\in{\mathbb{Z}}^{2}_{0}}\in\ell^{2}$ and
@@
Other possible choices of the norm in the cut-off function are possible for the 2D NSEs in vorticity formulation (1.1). However, the above cut-off has the highest smoothness that is compatible with Meyers’ estimate in Theorem A.1. Moreover, the choice of $H^{r}$ with $r>0$ will be crucial in Section 5 below, where we deal with 2D NSEs in velocity formulation, which are critical in $L^{2}$.
@@
Let $p\in[2,\infty)$. Fix $\nu>0$, $\theta^{\nu}\in\ell^{2}$ and $\zeta_{0}\in B^{1-2/p}_{2,p}(\mathbb{T}^{2})$. Let $\tau^{\nu}:\Omega\to[0,\infty]$ and $\zeta_{{\rm cut}}^{\nu}:[0,\tau^{\nu})\times\Omega\to H^{1}(\mathbb{T}^{2})$ be a stopping time and a progressive measurable process, respectively.
  • We say that $(\zeta_{{\rm cut}}^{\nu},\tau^{\nu})$ is a local $p$-solution to (4.1) if there exists a sequence of stopping times $(\tau^{\nu}_{n})_{n\geq 1}$ such that $\tau^{\nu}_{n}\leq\tau^{\nu}$ a.s. for all $n\geq 1$ and $\lim_{n\to\infty}\tau^{\nu}_{n}=\tau^{\nu}$ a.s., for which the following conditions hold for all $n\geq 1$:
    • $\zeta_{{\rm cut}}^{\nu}\in L^{p}(0,\tau^{\nu}_{n};H^{1}(\mathbb{T}^{2}))\cap C([0,\tau^{\nu}_{n}];B^{1-2/p}_{2,p}(\mathbb{T}^{2}))$ a.s.;
    • ${\mathcal{K}}\,\zeta_{{\rm cut}}^{\nu}\zeta_{{\rm cut}}^{\nu}\in L^{p}(0,\tau^{\nu}_{n};L^{2}(\mathbb{T}^{2};{\mathbb{R}}^{2}))$ a.s.;
    • a.s. for all $t\in[0,\tau^{\nu}_{n}]$ it holds that
@@
      $+\sqrt{2\mu}\int_{0}^{t}\mathbf{1}_{[0,\tau^{\nu}_{n}]}\Big((\theta^{\nu}_{k}\,\sigma_{k}\cdot\nabla)\zeta_{{\rm cut}}^{\nu}(s)\Big)_{k}\,\mathrm{d}\mathcal{W}_{\ell^{2}}.$
  • A local $p$-solution $(\zeta_{{\rm cut}}^{\nu},\tau^{\nu})$ to (4.1) is said to be a unique maximal (local) $p$-solution to (4.1) if for any other local $p$-solution $(\xi_{{\rm cut}}^{\nu},\lambda^{\nu})$ we have $\lambda^{\nu}\leq\tau^{\nu}$ a.s. and $\xi_{{\rm cut}}^{\nu}=\zeta_{{\rm cut}}^{\nu}$ a.e. on $0,\lambda^{\nu})\times\Omega$.
  • A unique maximal $p$-solution $(\zeta_{{\rm cut}}^{\nu},\tau^{\nu})$ to (4.1) is said to be a unique global $p$-solution to (4.1) if $\tau^{\nu}=\infty$ a.s.
In the above, $\mathcal{W}_{\ell^{2}}$ is as in (2.5). For the optimality of the regularity of the initial data, the reader is referred to the comments below Theorem A.1. Similar to Definition 2.2, if $p=2$, then we simply write ‘solution’ instead of ‘$p$-solution’.
@@
Fix $\nu\in(0,1)$ and $\mu>0$. Let $(\theta^{n})_{n\geq 1}\subseteq\ell^{2}$ be a sequence of normalized radially symmetric coefficients (i.e., satisfying (2.1)) such that
$\lim_{n\to\infty}\|\theta^{n}\|_{\ell^{\infty}}=0.$
Then there exists $p_{0}(\nu,\mu)>2$ for which the following assertion holds. If for $p\in(2,p_{0}]$, $r\in(0,1-2/p)$ and $R>0$ we have
@@
    there exists a unique solution $\zeta_{{\rm cut,d}}^{\nu}\in C([0,1];H^{r})\cap L^{2}(0,1;H^{1})$ on $[0,1]$ to
    (4.2) $\left\{\begin{aligned} \partial_{t}\zeta_{{\rm cut,d}}^{\nu}+\phi_{R,r}(\zeta_{{\rm cut,d}}^{\nu})\,[({\mathcal{K}}\,\zeta_{{\rm cut,d}}^{\nu}\cdot\nabla)\zeta_{{\rm cut,d}}^{\nu}]&=(\nu+\mu)\Delta\zeta_{{\rm cut,d}}^{\nu}&\text{ on }&\mathbb{T}^{2},\\ \zeta_{{\rm cut,d}}^{\nu}(0,\cdot)&=\zeta_{0}&\text{ on }&\mathbb{T}^{2};\end{aligned}\right$.
@@
(4.3) $\lim_{n\to\infty}\mathbf{P}\Big(\sup_{t\in[0,1]}\|\zeta_{{\rm cut},n}^{\nu}(t)-\zeta_{{\rm cut,d}}^{\nu}(t)\|_{H^{r_{0}}}\geq\varepsilon\Big)=0\ \text{ for all }\varepsilon>0,$
where $\zeta_{{\rm cut},n}^{\nu}$ is the unique global $p$-solution to (4.1) with $\theta=\theta^{n}$ and initial data $\zeta_{0,n}$.
As for Proposition 3.1, the key point is that the Sobolev space over $\mathbb{T}^{2}$ in which the convergence in (4.3) takes place has positive smoothness. Moreover, repeating the argument in Remark 3.4, one sees that $\nabla\zeta_{{\rm cut},n}^{\nu}\not\to\nabla\zeta_{{\rm cut,d}}^{\nu}$ as $n\to\infty$ in probability in $L^{2}(0,1;L^{2})$ in case $\zeta_{0}\not\equiv 0$. Moreover, in contrast to the linear case, the uniqueness of (weak) solutions to (4.2) is not immediate, and therefore we include it as an assumption.
The existence of a unique global $p$-solution to (4.1) is proven in Lemma 4.3 below. (Local) solutions to (4.2) on $[0,1]$ are as in Definition 4.1 with trivial noise, $p=2$ and life-time $\geq 1$. Note that the condition $\zeta_{{\rm det}}^{\nu}\in C([0,1];H^{r})$ is not a-priori given in the requirements for being local solutions, this is why it has been displayed in (2). Similarly, $\zeta_{{\rm det}}^{\nu}\in L^{2}(0,1;H^{1})$ is needed as life-time $\geq 1$ does not imply that $t\mapsto\|\zeta_{{\rm det}}^{\nu}(t)\|_{L^{2}}^{2}$ is integrable near $t=1$. Let us emphasize that uniqueness for (4.2) is only required among all maps on $[0,1]$ which belong to $C([0,1];H^{r})\cap L^{2}(0,1;H^{1})$.
The proof of Proposition 4.2 is given at the end of this subsection. Following the argument of Subsection 3.1, the key ingredients are the following estimates, which are uniform in the class of normalized radially symmetric $\theta^{\nu}$.
@@
As in the proof of Lemma 3.2, the key point is the independence of $(p_{0},C_{0})$ on $\theta$. As in the latter result, the key tools are the stochastic Meyers’ estimates of Theorem A.1. To handle the nonlinearity, we use the sub-criticality of $H^{r}$ for all $r>0$ in the case of 2D NSEs, both in vorticity and velocity form (see Section 5 below for the latter). The sub-criticality of $H^{r}$ is exploited via the following inequality (cf. [1, Lemmas 4.3 and 4.5]): For all $r<\frac{1}{2}$,
(4.5) $\big\|\nabla\cdot([{\mathcal{K}}\,\zeta]\,\zeta)\big\|_{H^{-1}}\lesssim\|\zeta\|_{L^{4}}^{2}\lesssim\|\zeta\|_{H^{1/2}}^{2}\lesssim\|\zeta\|_{H^{r}}^{1+\kappa_{r}}\|\zeta\|_{H^{1}}^{1-\kappa_{r}}$
where $\kappa_{r}=\frac{r}{1-r}>0$. For later use (see Section 5 below), let us note that the (4.5) also holds if ${\mathcal{K}}\,\zeta]\zeta$ is replaced by $v\otimes v$ with $v\in H^{1}$.
The above follows from the Sobolev embedding $H^{1/2}(\mathbb{T}^{2})\hookrightarrow L^{4}(\mathbb{T}^{2})$ as well as standard interpolation inequality. The sub-criticality of $H^{r}$ is encoded in the power $1-\kappa_{r}<1$ for the $H^{1}$-norm on the RHS(4.5). This fact will allow us to absorb the corresponding term via Young’s inequality.
Proof of Lemma 4.3.
@@
The proof is now split into three steps. We first prove the existence of a unique maximal $p$-solution $(\zeta_{{\rm cut}},\tau)$ with a corresponding blow-up criterion. Secondly, we prove that a localized (in time) version of (4.4) holds for such a maximal solution. Finally, in the third step, we prove that $\tau=\infty$ a.s. (i.e., $(\zeta_{{\rm cut}},\tau)$ is global) and that (4.4) holds.
Step 1: There exists a unique maximal (unique) $p$-solution $(\zeta_{{\rm cut}},\tau)$ to (4.1), and for all $T\in(0,\infty)$, the following blow-up criterion holds
@@
To prove the existence of a unique maximal (local) $p$-solution to (4.1), we employ [5, Theorem 4.8] (or [11, Theorem 4.7]). To this end, we rewrite (4.1) as a stochastic evolution equation on $X_{0}\stackrel{{\scriptstyle{\rm def}}}{{=}}H^{-1}$:
@@
where $\mathcal{W}_{\ell^{2}}$ is as in (2.5), and for $\zeta_{{\rm cut}}\in X_{1}\stackrel{{\scriptstyle{\rm def}}}{{=}}H^{1}$,
$A\zeta_{{\rm cut}}=-(\nu+\mu)\Delta\zeta_{{\rm cut}}$ $,\qquad F(\zeta_{{\rm cut}})=\phi_{R,r}(\zeta_{{\rm cut}})\,\nabla\cdot([{\mathcal{K}}\,\zeta_{{\rm cut}}]\,\zeta_{{\rm cut}}),$
@@
Comparing Definition 4.1 with [5, Definition 4.4], one readily see that unique maximal $p$-solution to (4.1) are equivalent to maximal $L^{p}_{0}$-solution to (4.7) (see also [6, Remark 5.6]). Here, we are implicitly using the reformulation (2.6) of the noise in terms of real-valued Brownian motions. Set
(4.8) $X_{\beta}\stackrel{{\scriptstyle{\rm def}}}{{=}}[X_{0},X_{1}]_{\beta}=H^{-1+2\beta}\ \text{ for }\ \beta\in(0,1),$
where $\cdot,\cdot]_{\beta}$ denotes the complex interpolation functor, see e.g., [18, Chapter 4]. By [5, Theorem 4.8] or [11, Theorem 4.7], to show the existence of a unique maximal $p$-solution, it suffices to show the existence of $1-1/p<\beta\leq\varphi<1$ such that $\varphi+\beta<2-1/p$, and for all $\zeta,\zeta^{\prime}\in X_{1}$,
(4.9) $\|F(\zeta)-F(\zeta^{\prime})\|_{X_{0}}$ $\lesssim(\|\zeta\|_{X_{3/4}}+\|\zeta^{\prime}\|_{X_{3/4}})\|\zeta-\zeta^{\prime}\|_{X_{3/4}}$
@@
Next, we prove (4.9). Without loss of generality, we assume that $\|\zeta^{\prime}\|_{H^{r}}\leq\|\zeta\|_{H^{r}}$. Under the latter condition, we write $F(\zeta)-F(\zeta^{\prime})=I_{\zeta,\zeta^{\prime}}+J_{\zeta,\zeta^{\prime}}$ where
(4.10) $I_{\zeta,\zeta^{\prime}}$ $=\phi_{R,r}(\zeta)\,\nabla\cdot([{\mathcal{K}}\,\zeta]\,\zeta)-\nabla\cdot([{\mathcal{K}}\,\zeta^{\prime}]\zeta^{\prime}),$
$J_{\zeta,\zeta^{\prime}}$ $=\nabla\cdot([{\mathcal{K}}\,\zeta^{\prime}]\,\zeta^{\prime})\,\big(\phi_{R,r}(\zeta)-\phi_{R,r}(\zeta^{\prime})\big).$
The first term on the RHS(4.9) clearly follows by estimating $I_{\zeta,\zeta^{\prime}}$ and using $X_{3/4}=H^{1/2}(\mathbb{T}^{2})\hookrightarrow L^{4}(\mathbb{T}^{2})$. Arguing as in (4.5), by interpolaton, for all $r<\frac{1}{2}$, we have
(4.11) $\big\|\nabla\cdot([{\mathcal{K}}\,\zeta^{\prime}]\,\zeta^{\prime})\big\|_{H^{-1}}\lesssim\|\zeta^{\prime}\|_{H^{1/2}}^{2}\lesssim\|\zeta^{\prime}\|_{H^{r}}\|\zeta^{\prime}\|_{H^{1-r}}.$
We emphasize that the choice of the space $H^{1-r}$ for the interpolation in (4.11) is, to a certain extent, arbitrary. For our purposes, we only need the spaces appearing on the RHS(4.11) to be of lower order compared to $X_{1}=H^{1}$.
Now, coming back to (4.9), as are assuming $\|\zeta\|_{H^{r}}\geq\|\zeta^{\prime}\|_{H^{r}}$, it follows that
(4.12) $|\phi_{R,r}(\zeta)-\phi_{R,r}(\zeta^{\prime})|=0\ \ \text{ if }\ \ \|\zeta^{\prime}\|_{H^{r}}\geq 2R.$
In particular,
[…]
$\ \lesssim_{R}\|\zeta^{\prime}\|_{H^{1-r}}\,\|\zeta-\zeta^{\prime}\|_{H^{r}}.$
Thus, it follows from (4.8) that, for all $\beta>1-\frac{1}{p}$ and $\varphi\in[(1-\frac{r}{2})\vee\beta,1)$,
(4.13) $\|J_{\zeta,\zeta^{\prime}}\|_{H^{-1}}\lesssim\|\zeta^{\prime}\|_{X_{\varphi}}\,\|\zeta-\zeta^{\prime}\|_{X_{\beta}}.$
Hence, the condition $\beta+\varphi<2-\frac{1}{p}$ follows by choosing $\beta=1-\frac{1}{p}+\varepsilon$ and $\varphi=1-2\varepsilon$ with $\varepsilon$ sufficiently small depending only on $r$ and $p$. Hence, (4.9) follows from (4.13) and the comments below (4.10).
Finally, (4.6) follows from [6, Theorem 4.10(3)] applied to (4.7).
@@
where $\inf\varnothing\stackrel{{\scriptstyle{\rm def}}}{{=}}\tau\wedge T$. The Meyers’ estimates of Theorem A.1 and a standard localization argument (see e.g., [5, Proposition 3.12(b)]) ensure that existence of $C_{1}(\nu,\mu)>0$ such that, for all $m\geq 1$, $p\in[2,p_{0}]$ and $u_{0}\in B^{1-2/p}_{2,p}$,
@@
Hence, $\|\zeta_{{\rm cut}}(t)\|_{B^{1-2/p}_{2,p}}^{p}+\int_{0}^{\tau\wedge T}\|\zeta_{{\rm cut}}(s)\|^{p}_{H^{1}}\,\mathrm{d}s<\infty$ a.s. on $\{\tau<T\}$, and therefore $\tau\geq T$ a.s. by (4.6). The arbitrariness of $T$ implies $\tau=\infty$ a.s., as desired.
Step 3: Proof of (4.4). The estimate (4.4) follows from (4.14) and the fact that $\tau=\infty$ a.s. by Step 2. ∎
@@
and that for all $\|\zeta_{0}\|_{H^{\delta}}\leq N$ the unique global solution $\zeta^{\nu}$ to (1.1) satisfies
@@
(4.15) $\left\{\begin{aligned} \partial_{t}\zeta_{{\rm det}}^{\nu}+({\mathcal{K}}\,\zeta_{{\rm det}}^{\nu}\cdot\nabla)\zeta_{{\rm det}}^{\nu}&=(\nu+\mu)\Delta\zeta_{{\rm det}}^{\nu}&\text{ on }&\mathbb{T}^{2},\\ \zeta_{{\rm det}}^{\nu}(0,\cdot)&=\zeta_{0}&\text{ on }&\mathbb{T}^{2}.\end{aligned}\right$.